OpenStudy (kimberly_pr):

Help please! This time it's on a different question. Using a directrix of y = −2 and a focus of (2, 6), what quadratic function is created?

OpenStudy (kimberly_pr):

@Mehek14 @DogzCatz @Maria195

OpenStudy (kimberly_pr):

@tom982

OpenStudy (anonymous):

The distance from a point on the parabola to the focus must be the same as the distance from that point to the line y=-2

OpenStudy (anonymous):

Let (x,y) be a point on the parabola, the distance squared from it to the point (2.6) is \[ d^2= (x-2)^2 +(y-6)^2 \]

OpenStudy (anonymous):

How much is the distance from (x,y) to the line y=-2?

OpenStudy (anonymous):

\[ d^2=(y+2)^2 \]

OpenStudy (anonymous):

Equate, Simplify and you get the answer

OpenStudy (kimberly_pr):

Umm, I'm not completely sure. Would it be f(x)= -1/8 (x-2)^2 -2?

OpenStudy (kimberly_pr):

...Am I right?

OpenStudy (kimberly_pr):

Or wait... Is it -1/16 (x-2)^2 -2? Now I'm not sure :/

OpenStudy (kimberly_pr):

Please help, I'm kind of confused now. :/

OpenStudy (kimberly_pr):

And I have another question. This one's a lot easier but my mind just blanked on me. Which polynomial identity will prove that 16 = 25 − 9?

OpenStudy (kimberly_pr):

Are you working it out?...

OpenStudy (kimberly_pr):

Nevermind, I just went ahead and answered because I couldn't wait any longer. I got it wrong for all of you who might use this as future reference :(

OpenStudy (anonymous):

The equation of the parabola should be \[ y=\frac{1}{16} \left(x^2-4 x+36\right) \]

OpenStudy (anonymous):

It is a parabola

OpenStudy (whpalmer4):

 It is good to be able to visualize these problems before working the equations. You have a directrix of \(y=-2\) and a focus of \((2,6)\). The parabola passes between the directrix and the focus (it is the set of all points equidistant from directrix and focus), so with the focus above the directrix, we have a parabola which opens upward. General form of a parabola which opens upward is \[y = a(x-h)^2+k\]where \((h,k)\) is the vertex of the parabola. It can also be written as \[y = \frac{1}{4p}(x-h)^2+k\]where \(p\)is half the distance between directrix and focus. If you have two points on the parabola, the former may be easier; if you have the focus and the directrix, the latter is easier. Distance from focus to directrix is simply \[| 6-(-2)| = 8\]so \(p=8/2=4\) That means our equation is \[y=\frac{1}{4(4)}(x-h)^2+k=\frac{1}{16}(x-h)^2+k\] To find the vertex, remember that the vertex is the point which is closest to both the focus and the directrix, and thus the point halfway between them. Our vertex must be at \(x=2\) and \(y = -2 + p = -2+4 = 2\) or \((2,2)\). That gives us \(h = 2,\ k=2\) and our equation is \[y = \frac{1}{16}(x-2)^2+2\] Here's a graph showing the various points of interest:

OpenStudy (whpalmer4):

And for your other question: "Which polynomial identity will prove that 16 = 25 − 9" \[\sqrt{25} = 5\]\[\sqrt{9}=3\]You can express\[25-9\]as the difference of two squares: \[(x-a)(x+a)= x*x+a*x-a*x-a*2 = x^2-a^2\] if we set \[x=5\]and \[a=3\]we have \[(5+3)(5-3) = 5^2-3*5+3*5-3*3 = 5^2-3^2\]but we also know that \[(5+3)(5-3) = 8*2 = 16\]therefore\[5^2-3^2 = 16\]even if we couldn't calculate the values of \(5^2\) and \(3^2\) for some bizarre reason. @Kimberly_PR